OpenStudy (anonymous):

A quality control technician at Cyberdyne Systems Corporation is checking each machine in a row until he finds one that does not operate, but he will check no more than three machines, even if they all operate. Assume each machine operates independently and has a probability of operating of 0.9. What is the expected number of operating machines?

ganeshie8 (ganeshie8):

geometric distribution

ganeshie8 (ganeshie8):

i think, expected value = mean

OpenStudy (anonymous):

No actually it's np

OpenStudy (anonymous):

THis is confusing.

OpenStudy (anonymous):

Sorry it's 1/p .

OpenStudy (anonymous):

But if I do that I don't get the right answer.

OpenStudy (anonymous):

P= P(3 work) + P(2 work, 1 fails) + P(1 works, 2 fails) + P(0 works, all 3 fail) So P=0.82

OpenStudy (anonymous):

1/p = 1.21 but the answer is 2.439

OpenStudy (anonymous):

Interesting 2/0.82 = 2.439 but that makes no sense.

OpenStudy (anonymous):

Are you saying it doesn't make sense because it's a decimal?

OpenStudy (anonymous):

No no.

OpenStudy (anonymous):

I'm saying it dsen't make sense because of that 2.

OpenStudy (anonymous):

Besides for geometric distribution E(x) =1/p . That 2 dosent even come into play here.

OpenStudy (anonymous):

And it should be geometric because it says until the first failure.

OpenStudy (anonymous):

Have you tried using a calculator? A geometric distribution one, I mean.

OpenStudy (anonymous):

We aren't allowed to use anything more sophisticated than a hand calculator.

OpenStudy (anonymous):

Arragh. THis is frustrating me >.> .

OpenStudy (anonymous):

A single mutiple choice should not be so hard >.< .

OpenStudy (anonymous):

Well, the expected value would have to be between 0 and 1 at least.

OpenStudy (anonymous):

No it has be be between 0 and 3.

OpenStudy (anonymous):

SInce he checks 3 machines.

OpenStudy (anonymous):

The probability that one fails is: \[ (1-0.9)+(0.9)^1(1-0.9) + (0.9)^2(1-0.9) \]

OpenStudy (anonymous):

\[ =0.271 \]

OpenStudy (anonymous):

Oh wait, I see I misread it.

OpenStudy (anonymous):

Regardless, we can still use \(1-0.271\) for the probability that none fail. This means \[ 0\times (1-0.9)+1\times (0.9)^1(1-0.9) + 2\times (0.9)^2(1-0.9) + 3\times (1-0.271) \]

OpenStudy (kirbykirby):

I suggest using a table of values like X | ----------- f(x) |

OpenStudy (anonymous):

\[ =2.439 \]

OpenStudy (anonymous):

Dido, do you have an answer to check it?

OpenStudy (anonymous):

Yeah that's right!

OpenStudy (anonymous):

Well, it is hard to understand what I did?

OpenStudy (anonymous):

A little.

OpenStudy (anonymous):

The fallacy here is that the distribution at \(\Pr(X=3)\) is not geometric.

OpenStudy (anonymous):

Why not? It says until the first one fails.

OpenStudy (anonymous):

Doing so would be assuming that \(\Pr(X=4) \geq 0\). But how can that be possible if he would never check 4 of them?

OpenStudy (anonymous):

Whoops, meant to pus \(\Pr(X=4)\gt 0\).

OpenStudy (anonymous):

AHh right that makes a lot of sense yes.

OpenStudy (anonymous):

Now explain what you did. It's P(3 fail)+P(2 fail)+P(1 fail)+P(0 fail) right?

OpenStudy (anonymous):

I used geometric to calculate \(\Pr(X=k)\) when \(k=0,1,2\)

OpenStudy (anonymous):

For \(\Pr(X=3)\), I used total probability: \[ 1 = \Pr(X=0)+ \Pr(X=1)+ \Pr(X=2)+ \Pr(X=3) \]

OpenStudy (anonymous):

Once you calculate the probabilities, getting the mean shouldn't be an issue.

OpenStudy (anonymous):

I know. It was just difficult to see this.

OpenStudy (anonymous):

\[ \mu =\sum_{k=0}^{3}k\times \Pr(X=k) \]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

If you're still there @wio , why is the second term in your equation does not have a squared sign?

OpenStudy (anonymous):

First term would have \((1-0.9)^0\), but I figured you wouldn't need it.

OpenStudy (anonymous):

I mean would have \((0.9)^0\).

OpenStudy (anonymous):

For 0 operating, 1 fail. Second term is 1 operating 1 fail.

OpenStudy (anonymous):

I had: \[0(0.271)^3+1(0.90)(0.10)^2+2(0.90)^2(0.10)+3(0.90)^3\]

OpenStudy (anonymous):

Should it not be 1 operation 2 fail?

OpenStudy (anonymous):

operating*

OpenStudy (anonymous):

Actually I guess it makes sense since he only needs 1 failure. SO we don't square it.

OpenStudy (anonymous):

Right? @wio

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

Thanks :P .